OpenStudy (anonymous):

Estimate the length of the radius if the area of the circle is 54 ft2. A. 17.19 ft B. 4.15 ft C. 8.29 ft D. 8.60 ft

OpenStudy (anonymous):

It's B

OpenStudy (anonymous):

A = pi r^2 -> r = sqrt ( A/ pi) = sqrt ( 54/ pi) = 17.19 ft

OpenStudy (anonymous):

is it A or b?

OpenStudy (anonymous):

It's B.

OpenStudy (anonymous):

@chlorophyll sayd its A

OpenStudy (anonymous):

You should show your reason, atjari :)

OpenStudy (anonymous):

o.O now its a b or c

sam (.sam.):

\[\pi(r)^{2}\]

sam (.sam.):

r=4.415

sam (.sam.):

B

OpenStudy (anonymous):

its B?

sam (.sam.):

thought he wait to find diameter, lol

sam (.sam.):

yes its B

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

chlorophyll I thnk u forgot to find the sqrt.

OpenStudy (anonymous):

I'm wrong, I missed the sqrt ( ) --> 4.15 => B

OpenStudy (mertsj):